The modulus of rigidity and Poisson’s ratio of the material of a wire are 2.87 × 1010 N/m2, and 0.379 respectively. Estimate the Young’s modulus of the material of the wire. 


1
Expert's answer
2021-03-12T09:22:36-0500

The relation between the modulus of rigidity "G", Poisson’s ratio "\nu", and the Young’s modulus "E" is the following:


"E = 2G(1+\nu)"

Substituting the numbers, obtain:


"E = 2\cdot 2.87\times 10^{10}\cdot (1 + 0.379) \approx 7.92\times 10^{10}\dfrac{N}{m^2}"

Answer. "7.92\times 10^{10}\dfrac{N}{m^2}".